OpenStudy (anonymous):

Determine whether each equation has the given ordered pair as a solution. 5x − 3y = 15; (0, 5)

OpenStudy (anonymous):

Let x=0 5x − 3y = 15 5(0) − 3y = 15 0-3y=15 3y=15 -- --- 3 3 y=5 (0,5)

OpenStudy (anonymous):

thank you cause i wasnt sure of the question

OpenStudy (amistre64):

?? you got an error there

OpenStudy (anonymous):

can you give me a medal?

OpenStudy (amistre64):

for starters; when x = 0 there is nothing over there and your left with -3y = 15 -3(5) does not equal 15

OpenStudy (anonymous):

amistre where would be the error ok hold on xok

OpenStudy (anonymous):

-15

OpenStudy (anonymous):

whoops

OpenStudy (amistre64):

xok dropped the - sign from the y

OpenStudy (amistre64):

but its good to know how to dbl chk to spot for those ..... the steps were good, just missed it by thaaaaaatt much lol

OpenStudy (anonymous):

so -3/15 y=-5?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

okay thax for checking amistre

OpenStudy (amistre64):

just plug in your point values and see if it matches really (x=0, y=5); 5(0)-3(5)=15 0 -15 = 15 -15 = 15 is a contradiction; so the point aint part of the equation

OpenStudy (anonymous):

can we have medals?

OpenStudy (amistre64):

since ya tried :)

OpenStudy (anonymous):

grrr :)

OpenStudy (anonymous):

so amistre im safe with ur steps?

OpenStudy (anonymous):

:(

OpenStudy (amistre64):

yep; if the point fits and makes it true; its good .... otherwise toss it

OpenStudy (anonymous):

thanx both no sad faces u guys are helping me ! i appreciate it ok

OpenStudy (amistre64):

:) youre welcome

OpenStudy (anonymous):

It's not a solution. The answer would come out as -15 instead of 15

OpenStudy (anonymous):

the solution is if there are 2 lines... the solution would be where the two lines meet